OpenStudy (anonymous):

Can You walk me through this? Find the solution of 3times square root of x plus 6 , equals negative 12

OpenStudy (lucaz):

\[\large{3\times \sqrt{x-6}=-12}\]

OpenStudy (lucaz):

@iforgothowtomath this is the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[3(\sqrt{x+6})=-12\]

OpenStudy (anonymous):

well the 3 is just in front of the square root

OpenStudy (anonymous):

Divide by 3

OpenStudy (anonymous):

\[\sqrt{x+6}=-4\]

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

ok so how do i gte it out of square root

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x+6=16\] cuz u square both sides.

OpenStudy (anonymous):

x=10 after you subtract 6 from both sides. :)

OpenStudy (anonymous):

oh okay i did that

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

oh cool ok i did that

OpenStudy (anonymous):

ok so is it extraneous because it has a solution

OpenStudy (anonymous):

and thank you for your help